# Benchmarking Ultra-Low-Power NPUs

### v2

@@ 4.7. Limitations
Several limitations should be considered when interpreting our benchmarking results:
Frequency Standardization: While enforcing a uniform CPU frequency across all platforms enables direct comparison of architectural efficiencies, it fails to showcase each platform’s peak performance – many of the benchmark platforms can operate at higher frequencies than evaluated.
Fixed Quantization Bit-Width: Our standardized INT8 quantization approach, while enabling fair comparison, does not leverage the full capabilities of platforms supporting lower bit-width operations (, MAX78000’s 1/2/4-bit support) or higher precision computation (, HX-WE2’s
FLOAT16/FLOAT32 support). We also only focus on quantization as a means of reducing model size, excluding other optimization methods.
@@
Operator Support: Similarl, by ensuring all models are fully NPU-compatible across the various evaluated platforms, we negate the impact of unsupported NN operators. Further work should examine performance scaling across platforms with different sets of supported operators, using more complex or unmodified models, alongside precision-optimized models for each platform, and the impact of platform-specific architectural optimizations.
Development Toolchain Maturity We focus solely on performance metrics in this study. However, another often overlooked factor when selecting a target platform, deserving attention in future studies, is the relative maturity of its development ecosystem and model optimization tools.
@@ 5. RELATED WORK
Benchmarking NN Models on Constrained Hardware: A growing body of literature has explored NN benchmarking on constrained and mobile computing platforms. Japana et al.’s MLPerf benchmark introduced the first industry-standard open-source framework for performance evaluation of NNs on mobile devices equipped with diverse NN accelerators and software stacks (janapa2022mlperf, ). Laskaridis et al. recently investigated the efficiency of large language models (LLMs) on various SOTA mobile platforms, including Android, iOS and Nvidia Jetson devices (laskaridis2024melting, ). Reuther et al. explored the performance and power characteristics of a wide range of NN accelerators, spanning cellular GPUs, FPGA accelerators, up to data center hardware (reuther2019survey, ). However, previous work on NPU platforms has been limited to application-level performance assessments (wang2022bed, ; song2024tada, ) or single-platform standalone benchmarks (moss2022ultra, ; huang2024energy, ). Furthermore, existing single-platform benchmarks often overlook certain operations in the end-to-end model pipeline (clay2022benchmarking, ). Hence, to the best of our knowledge, our work details the first side-by-side and fine-grained benchmarking study of NN models across a number of commercially-available NPU platforms.
@@
Efficient On-Device Inference: Deploying NNs on MCUs is constrained by the underlying hardware’s memory capacity and throughput, with power consumption also often emerging as a bottleneck (afzal2022battery, ; zhao2022towards, ). Numerous works have explored model compression (choudhary2020comprehensive, ; lin2020mcunet, ; zawish2024complexity, ), the design of more efficient NN operators and architectures for lower resource usage (pan2023reusing, ; tarnawski2020efficient, ; li2017simple, ), and adaptive NN inference based on input complexity and workload (ee, ; delight, ; hydranets, ). Various hardware-based optimizations have also been studied, such as parallel dataflow processing (gong2025dex, ). Our work aims to further advance efficient NN deployment across NPU platforms by identifying the current SOTA alongside existing bottlenecks.
@@ 6. CONCLUSION
Our comprehensive evaluation of various NN models across commercially-available NPUs reveals both expected trends as well as unexpected findings, contributing to the growing body of knowledge on embedded neural computation.
The significant performance advantages of dedicated neural acceleration are clearly demonstrated, with specialized platforms achieving up to two orders of magnitude higher energy-efficiency compared to general-purpose MCUs. We also highlight that theoretical computational capacity (GOPs) alone is an insufficient predictor of real-world performance. The stage-by-stage breakdown of model inference reveals critical bottlenecks on certain platforms – particularly in memory I/O operations – alongside key insights for future work in hardware and model design. We urge developers to consider trade-offs in latency, energy-efficiency, model complexity, and operational flexibility to achieve optimal performance in real-world deployments. We open-source our benchmarking framework and hope its use can streamline cross-platform model compilation and evaluation.
@@ 7. ACKNOWLEDGMENTS
90 Model Stage MAX78k (C MAX78k (R) GAP8 NXP-MCXN947 HX-WE2 (S) HX-WE2 (P) STM32H7A3ZI ESP32 Time Power Time Power Time Power Time Power Time Power Time Power Time Power Time Power NAS Initialization 0.07±0.001 13.67±0.02 0.17±0.002 12.84±0.03 12.94±0.237 102.93±3.11 0.24±0.001 25.40±3.39 2.60±0.002 91.50±0.22 2.60±0.001 91.44±0.19 0.43±0.005 47.38±0.07 86.73±0.001 105.93±0.09 Memory I/O 21.24±0.004 25.05±0.05 35.53±0.003 22.66±0.03 2.66±0.007 38.58±0.95 0.83±0.001 24.52±0.84 0.12±0.001 97.55±1.25 0.12±0.001 97.01±0.71 Inference 4.63±0.002 80.41±0.08 4.64±0.002 79.03±0.05 77.40±0.470 117.22±1.47 24.86±0.002 36.38±0.65 8.99±0.022 112.35±0.76 8.32±0.008 114.62±0.76 550.01±0.012 54.91±0.09 468.38±0.001 151.99±0.31 Post-Processing 0.02±0.001 25.66±0.04 0.11±0.001 22.43±0.06 0.08±0.001 33.23±1.28 0.02±0.001 24.03±1.19 0.01±0.001 91.19±1.31 0.01±0.001 91.17±1.39 0.02±0.001 46.01±0.05 0.02±0.001 101.93±1.86 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.53±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 ResNet Initialization 0.07±0.001 13.87±0.05 0.17±0.001 12.82±0.07 10.13±0.001 145.90±3.59 0.24±0.001 25.14±1.99 2.59±0.001 94.27±0.23 2.60±0.001 93.51±2.01 0.43±0.002 48.32±0.07 87.49±0.059 106.41±0.07 Memory I/O 26.53±0.002 25.14±0.05 44.87±0.003 22.64±0.04 2.47±0.110 38.39±0.31 0.91±0.001 24.89±0.99 0.12±0.001 99.15±0.45 0.14±0.001 97.74±1.37 Inference 2.89±0.001 47.12±0.09 2.96±0.001 46.86±0.07 38.45±0.081 113.47±0.40 12.97±0.002 36.69±0.52 8.30±0.001 110.13±0.79 8.54±0.015 109.71±2.35 282.00±0.010 54.98±0.08 318.15±0.001 144.91±0.27 Post-Processing 0.12±0.001 26.10±0.06 0.96±0.002 22.45±0.05 0.04±0.001 37.41±1.55 0.04±0.001 24.60±1.95 0.04±0.001 93.19±0.15 0.04±0.001 91.95±1.27 0.08±0.002 50.47±0.09 0.05±0.001 102.79±3.71 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.62±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 SimpleNet Initialization 0.07±0.001 13.71±0.04 0.17±0.002 12.81±0.03 10.02±0.002 137.44±3.95 0.24±0.001 25.23±0.42 2.56±0.003 95.75±0.49 2.60±0.001 91.51±0.21 0.43±0.001 48.13±0.04 87.37±0.001 106.58±0.66 Memory I/O 12.14±0.003 24.46±0.07 20.22±0.003 22.27±0.04 2.60±0.076 38.51±0.69 0.91±0.001 24.84±1.69 0.12±0.001 91.11±1.77 0.12±0.001 96.97±0.69 Inference 2.68±0.002 50.89±0.04 2.67±0.003 49.86±0.05 38.23±0.117 122.12±3.70 12.49±0.001 35.61±0.81 8.10±0.001 107.15±0.69 7.13±0.001 106.84±0.77 276.03±0.011 55.19±0.06 283.21±0.002 139.96±1.84 Post-Processing 0.11±0.001 25.62±0.07 0.86±0.001 22.51±0.05 0.04±0.001 36.83±0.39 0.14±0.001 24.87±0.21 0.04±0.001 90.54±0.14 0.04±0.001 90.47±0.21 0.10±0.002 46.10±0.07 0.05±0.001 105.11±3.25 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.73±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 Autoenc Initialization 0.07±0.001 13.75±0.06 0.17±0.001 12.86±0.05 8.17±0.076 89.22±1.28 0.23±0.001 24.95±0.20 2.60±0.002 90.79±0.70 2.60±0.001 96.04±1.01 0.12±0.002 48.17±0.06 19.76±0.002 103.51±0.59 Memory I/O 9.66±0.001 25.04±0.08 15.40±0.003 22.63±0.06 2.63±0.042 38.55±0.56 0.22±0.001 24.01±1.21 0.03±0.001 92.75±1.83 0.03±0.001 94.03±0.62 Inference 0.14±0.001 81.67±0.09 0.16±0.001 81.34±0.07 0.67±0.015 84.35±1.25 0.51±0.003 31.49±0.44 0.45±0.001 92.40±1.19 0.75±0.001 99.02±1.56 5.01±0.110 56.11±0.09 7.11±0.001 157.17±0.17 Post-Processing - - - - - - - - - - - - - - - - Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.53±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 YoloV1 Initialization 0.07±0.001 13.90±0.03 0.17±0.001 12.87±0.02 7.46±0.001 96.28±2.90 0.28±0.001 25.66±0.39 2.60±0.003 100.94±1.01 2.60±0.003 100.45±0.59 3.64±0.002 49.04±0.05 628.97±0.002 106.49±0.163 Memory I/O 8.84±0.003 41.82±0.04 11.99±0.002 35.55±0.04 2.83±0.037 38.10±0.11 0.77±0.001 25.41±1.61 1.11±0.001 100.31±3.09 1.11±0.001 101.95±2.90 Inference 2.61±0.002 21.47±0.04 2.62±0.001 21.13±0.05 36.11±0.047 113.67±2.39 17.18±0.001 28.43±1.66 8.48±0.001 112.63±1.30 7.30±0.001 111.69±0.97 336.00±0.013 54.99±0.08 536.22±0.001 129.74±0.107 Post-Processing 0.54±0.001 25.87±0.06 3.82±0.001 22.41±0.04 1.82±0.002 39.89±0.62 1.27±0.001 24.08±0.15 0.42±0.001 93.63±1.45 0.42±0.001 95.65±0.79 0.48±0.001 45.71±0.04 0.53±0.001 109.57±0.08 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 22.91±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.74±0.07

### v3

@@ 4.7. Limitations
Frequency Standardization: While enforcing a uniform CPU frequency across all platforms enables direct comparison of architectural efficiencies, it fails to showcase each platform’s peak performance – for example, many of the benchmark platforms can operate at higher frequencies than evaluated. Many platforms also combine DVFS with selective domain gating, so the net benefit of frequency scaling depends on both the workload and the platform.
Fixed Quantization Bit-Width: While the fixed-INT8 quantization approach enables cross-platform fairness, it also masks important trade-offs. Certain NPUs, such as the MAX78000, can gain substantial latency and energy savings from ultra-low bit-width quantization (1–4 bits), whereas others, like the HX-WE2, can preserve accuracy for sensitive workloads by exploiting FLOAT16 or FLOAT32 acceleration.
@@
Operator Support: Similarly, by ensuring all models are fully NPU-compatible across the various evaluated platforms, we negate the impact of unsupported NN operators. Further work should examine performance scaling across platforms with different sets of supported operators, using more complex or unmodified models, alongside precision-optimized models for each platform, and the impact of platform-specific architectural optimizations.
@@ 5. RELATED WORK
Benchmarking NN Models on Constrained Hardware: Japana et al.’s MLPerf benchmark introduced the first industry-standard open-source framework for performance evaluation of NNs on mobile devices equipped with diverse NN accelerators and software stacks (janapa2022mlperf, ). Laskaridis et al. recently investigated the efficiency of large language models (LLMs) on various SOTA mobile platforms, including Android, iOS and Nvidia Jetson devices (laskaridis2024melting, ). Reuther et al. explored the performance and power characteristics of a wide range of NN accelerators, spanning cellular GPUs, FPGA accelerators, up to data center hardware (reuther2019survey, ). However, existing work on NPU platforms has been limited to application-level performance assessments (wang2022bed, ; song2024tada, ) or single-platform standalone benchmarks (moss2022ultra, ; huang2024energy, ).
@@
Efficient On-Device Inference: Numerous works have explored model compression (choudhary2020comprehensive, ; lin2020mcunet, ; zawish2024complexity, ), the design of more efficient NN operators/architectures for lower resource usage (pan2023reusing, ; tarnawski2020efficient, ; li2017simple, ), and adaptive NN inference based on input complexity and workload (ee, ; delight, ; hydranets, ). Various hardware-based optimizations have also been studied, such as parallel dataflow processing (gong2025dex, ). Our work aims to further advance efficient NN deployment across NPU platforms by identifying current hardware bottlenecks.
@@ 6. CONCLUSION
Our evaluation of NN models across commercially-available NPUs reveals both expected trends and surprising findings. We show that dedicated neural accelerators deliver up to two orders of magnitude higher energy-efficiency than MCUs, but that theoretical capacity (, GOPs) alone poorly predicts real-world performance. Our stage-by-stage breakdown reveals critical bottlenecks on certain platforms – particularly in memory I/O operations – alongside key insights for future work in hardware and model design. We encourage developers to consider trade-offs in latency, energy-efficiency, model complexity, and flexibility for optimal deployment. We open-source our benchmarking toolchain and hope its use can streamline cross-platform compilation and evaluation.
@@ 7. ACKNOWLEDGMENTS
| Model | Stage | MAX78k (C) | MAX78k (R) | GAP8 | NXP-MCXN947 | HX-WE2 (S) | HX-WE2 (P) | STM32H7A3ZI | ESP32 | ||||||||
| Time | Power | Time | Power | Time | Power | Time | Power | Time | Power | Time | Power | Time | Power | Time | Power | ||
| NAS | Initialization | 0.074±0.001 | 13.67±0.02 | 0.169±0.002 | 12.84±0.03 | 12.939±0.237 | 102.93±3.11 | 0.219±0.001 | 112.29±3.39 | 2.597±0.002 | 91.50±0.22 | 2.601±0.001 | 91.44±0.19 | 0.429±0.005 | 47.38±0.07 | 86.727±0.001 | 105.93±0.09 |
| Memory I/O | 21.241±0.004 | 25.05±0.05 | 35.529±0.003 | 22.66±0.03 | 2.661±0.007 | 38.58±0.95 | 0.049±0.001 | 117.26±0.84 | 0.123±0.001 | 97.55±1.25 | 0.123±0.001 | 97.01±0.71 | — | — | — | — | |
| Inference | 4.625±0.002 | 80.41±0.08 | 4.642±0.002 | 79.03±0.05 | 77.396±0.470 | 117.22±1.47 | 27.267±0.002 | 118.03±0.65 | 8.988±0.022 | 112.35±0.76 | 8.319±0.008 | 114.62±0.76 | 550.010±0.012 | 54.91±0.09 | 468.381±0.001 | 151.99±0.31 | |
| Post-Processing | 0.016±0.001 | 25.66±0.04 | 0.111±0.001 | 22.43±0.06 | 0.076±0.001 | 33.23±1.28 | 0.010±0.001 | 109.62±1.19 | 0.007±0.001 | 91.19±1.31 | 0.007±0.001 | 91.17±1.39 | 0.015±0.001 | 46.01±0.05 | 0.017±0.001 | 101.93±1.86 | |
| Idle | — | 13.21±0.02 | — | 10.87±0.01 | — | 33.67±0.35 | — | 105.71±0.14 | — | 89.09±1.30 | — | 89.09±1.30 | — | 38.13±0.05 | — | 77.73±0.07 | |
| ResNet | Initialization | 0.074±0.001 | 13.87±0.05 | 0.169±0.001 | 12.82±0.07 | 10.133±0.001 | 145.90±3.59 | 0.221±0.001 | 110.37±1.99 | 2.593±0.001 | 94.27±0.23 | 2.602±0.001 | 93.51±2.01 | 0.432±0.002 | 48.32±0.07 | 87.489±0.059 | 106.41±0.07 |
| Memory I/O | 26.526±0.002 | 25.14±0.05 | 44.868±0.003 | 22.64±0.04 | 2.471±0.110 | 38.39±0.31 | 0.050±0.001 | 115.95±0.99 | 0.123±0.001 | 99.15±0.45 | 0.136±0.001 | 97.74±1.37 | — | — | — | — | |
| Inference | 2.893±0.001 | 47.12±0.09 | 2.964±0.001 | 46.86±0.07 | 38.447±0.081 | 113.47±0.40 | 16.067±0.002 | 118.57±0.52 | 8.295±0.001 | 110.13±0.79 | 8.535±0.015 | 109.71±2.35 | 282.000±0.010 | 54.98±0.08 | 318.149±0.001 | 144.91±0.27 | |
| Post-Processing | 0.123±0.001 | 26.10±0.06 | 0.963±0.002 | 22.45±0.05 | 0.037±0.001 | 37.41±1.55 | 0.039±0.001 | 109.55±1.95 | 0.038±0.001 | 93.19±0.15 | 0.038±0.001 | 91.95±1.27 | 0.082±0.002 | 50.47±0.09 | 0.052±0.001 | 102.79±3.71 | |
| Idle | — | 13.21±0.02 | — | 10.87±0.01 | — | 33.67±0.35 | — | 105.71±0.14 | — | 89.09±1.30 | — | 89.09±1.30 | — | 38.13±0.05 | — | 77.73±0.07 | |
| SimpleNet | Initialization | 0.074±0.001 | 13.71±0.04 | 0.168±0.002 | 12.81±0.03 | 10.017±0.002 | 137.44±3.95 | 0.221±0.001 | 108.98±0.42 | 2.560±0.003 | 95.75±0.49 | 2.597±0.001 | 91.51±0.21 | 0.431±0.001 | 48.13±0.04 | 87.372±0.001 | 106.58±0.66 |
| Memory I/O | 12.137±0.003 | 24.46±0.07 | 20.220±0.003 | 22.27±0.04 | 2.598±0.076 | 38.51±0.69 | 0.049±0.001 | 117.54±1.69 | 0.123±0.001 | 91.11±1.77 | 0.123±0.001 | 96.97±0.69 | — | — | — | — | |
| Inference | 2.657±0.002 | 50.89±0.04 | 2.670±0.003 | 49.86±0.05 | 38.226±0.117 | 122.12±3.70 | 13.523±0.001 | 118.40±0.81 | 8.099±0.001 | 107.15±0.69 | 7.134±0.001 | 106.84±0.77 | 276.000±0.011 | 55.19±0.06 | 283.210±0.002 | 139.96±1.84 | |
| Post-Processing | 0.114±0.001 | 25.62±0.07 | 0.864±0.001 | 22.51±0.05 | 0.037±0.001 | 36.83±0.39 | 0.038±0.001 | 109.33±0.21 | 0.039±0.001 | 90.54±0.14 | 0.038±0.001 | 90.47±0.21 | 0.102±0.002 | 46.10±0.07 | 0.053±0.001 | 105.11±3.25 | |
| Idle | — | 13.21±0.02 | — | 10.87±0.01 | — | 33.67±0.35 | — | 105.71±0.14 | — | 89.09±1.30 | — | 89.09±1.30 | — | 38.13±0.05 | — | 77.73±0.07 | |
| AI8XAutoEnc | Initialization | 0.074±0.001 | 13.75±0.06 | 0.168±0.001 | 12.86±0.05 | 8.174±0.076 | 89.22±1.28 | 0.279±0.001 | 109.14±0.20 | 2.599±0.002 | 90.79±0.70 | 2.601±0.001 | 96.04±1.01 | 0.124±0.002 | 48.17±0.06 | 19.758±0.002 | 103.51±0.59 |
| Memory I/O | 9.663±0.001 | 25.04±0.08 | 15.398±0.003 | 22.63±0.06 | 2.630±0.042 | 38.55±0.56 | 0.150±0.001 | 112.87±1.21 | 0.031±0.001 | 92.75±1.83 | 0.031±0.001 | 94.03±0.62 | — | — | — | — | |
| Inference | 0.143±0.001 | 81.67±0.09 | 0.156±0.001 | 81.34±0.07 | 0.696±0.015 | 84.35±1.25 | 19.246±0.003 | 112.19±0.44 | 0.451±0.001 | 92.40±1.19 | 0.749±0.001 | 99.02±1.56 | 5.010±0.110 | 56.11±0.09 | 7.106±0.001 | 157.17±0.17 | |
| Post-Processing | — | — | — | — | — | — | — | — | — | — | — | — | — | — | — | — | |
| Idle | — | 13.21±0.02 | — | 10.87±0.01 | — | 33.67±0.35 | — | 105.71±0.14 | — | 89.09±1.30 | — | 89.09±1.30 | — | 38.13±0.05 | — | 77.73±0.07 | |
| YoloV1 | Initialization | 0.074±0.001 | 13.90±0.03 | 0.169±0.001 | 12.87±0.02 | 7.457±0.001 | 96.28±2.90 | 0.277±0.001 | 108.85±0.39 | 2.599±0.003 | 100.94±1.01 | 2.598±0.003 | 100.45±0.59 | 3.639±0.002 | 49.04±0.05 | 628.970±0.002 | 106.49±0.163 |
| Memory I/O | 8.841±0.003 | 41.82±0.04 | 11.995±0.002 | 35.55±0.04 | 2.828±0.037 | 38.10±0.11 | 0.514±0.001 | 110.51±1.61 | 1.106±0.001 | 100.31±3.09 | 1.106±0.001 | 101.95±2.90 | — | — | — | — | |
| Inference | 2.612±0.002 | 21.47±0.04 | 2.623±0.001 | 21.13±0.05 | 36.109±0.047 | 113.67±2.39 | 19.251±0.001 | 113.31±1.66 | 8.477±0.001 | 112.63±1.30 | 7.295±0.001 | 111.69±0.97 | 336.000±0.013 | 54.99±0.08 | 536.222±0.001 | 129.74±0.107 | |
| Post-Processing | 0.538±0.001 | 25.87±0.06 | 3.818±0.001 | 22.41±0.04 | 1.817±0.002 | 39.89±0.62 | 0.112±0.001 | 109.35±0.15 | 0.419±0.001 | 93.63±1.45 | 0.420±0.001 | 95.65±0.79 | 0.476±0.001 | 45.71±0.04 | 0.534±0.001 | 109.57±0.08 | |
| Idle | — | 13.21±0.02 | — | 10.87±0.01 | — | 33.67±0.35 | — | 105.71±0.14 | — | 89.09±1.30 | — | 89.09±1.30 | — | 38.13±0.05 | — | 77.736±0.07 | |